# Distributional Regression with Tabular Foundation Models: Evaluating Probabilistic Predictions via Proper Scoring Rules

### 09 Mar 2026 v1

@@ 7 Outlook
While we discussed how to properly measure the goodness of a probabilistic forecast - namely by applying proper scoring rules, we highlight, that this discussion is not finished. Gneiting pointed out in Gneiting and Ranjan (2011) that situations with asymmetric risk (finance, weather forecasting etc.) may require to optimize for specific point estimates (via scoring functions suitable to elicit a specific point estimate). In such a case, we wonder how tabular foundation models would adapt to specific user supplied scoring functions (so that optimality is guaranteed). We point out a conflict of
  • a foundation model which is not altered by a user at all after pretraining is finished
  • the need to take into account that the optimal scoring function depends on the use-case at hand (must be supplied by the user since the user intention dictates the utility).
We see that Fine-Tuning of foundation models could be one route to enable optimal predictions for arbitrary scoring functions. Another route could be to enable task-token conditional probability prediction in case of asymmetric risk structures (a user would prompt the model with a desired scoring rule which should be optimal, but this behaviour would need to be pretrained). Such a task-token conditional probability prediction could output probability estimates which are optimal under some parametrizable family of scoring rules (like e.g. the beta energy score family). However there is infinitely many scoring rules and different families of scoring rules so that it is unlcear if pre-training with taks-tokens can in general adapt to arbitrary scoring rules as objectives. Potentially, tabular foundation models for high stakes applications need to make sensible choices for task conditioning for certain applications (finance, pharma production, etc.).
7.1 Example: Weighted CRPS (wCRPS)
Accounting for asymmetric risk structures, a modeler can choose to apply the different versions of weighted CRPSGneiting and Ranjan (2011) - a family of strictly proper scoring rules which enables e.g. emphasizing prediction quality for the left/right tail of a distribution which makes sense when dealing with e.g. energy forecasting or forecasting weather events with asymmetric cost.
7.2 Example: Interval Score
The Interval ScoreGneiting and Raftery (2007) targets the $(1-\alpha)$ central interval and is strictly proper, directly answering "how well is the interval estimated?" for a given significance level $\alpha$:
where $[l,u]$ are the $\alpha/2$ and $1-\alpha/2$ quantiles of the prediction $F$.
This specific scoring rule highlights that different scoring rules favor optimization in distinct regions of the probability mass function (PMF); in the extreme case of the interval score, the two boundary quantiles are of primary importance. Because we hypothesize that each scoring rule aligns with a specific utility, general-purpose models may fail to achieve optimality for specialized downstream tasks where the cost of error is non-uniform across the domain $\Omega$.
@@ 9 Appendix
9.1 Toy-Models to investigate finite sample effects
@@ 9.1.2 Toy Models
Probabilistic Classification Models These models predict probabilities across $N=50$ discrete bins of the $y$-variable. The neural networks are optimized using the Adam optimizer and are simple multi-layer perceptron (MLP) featuring an input layer, two hidden layers with 64 neurons each using $\tanh$ activations, and a Softmax output layer for $n$ bins.
@@
Loss Function RPS Loss CrossEntropyLoss
@@
  • PyTorch Ranked Probability Score (RPS) Model: A neural network trained directly to minimize the expected CRPS, which optimizes the accuracy of the cumulative distribution function (CDF).
  • XGBoost RPS Model: An XGBoost vectorleaf classifier utilizing a custom objective function that directly optimizes the Continuous Ranked Probability Score
@@
    7 Improved RPS Objective for XGBoost.
@@
    46def rps_metric(preds, dtrain):
    47 """Evaluation metric for RPS."""
    48 labels = dtrain.get_label().astype(int)
    49 num_samples = labels.shape[0]
    50 num_bins = preds.size // num_samples
    51
    52 preds = preds.reshape(num_samples, num_bins)
    53 probs = F.softmax(torch.tensor(preds), dim=1).numpy()
    54
    55 target_one_hot = np.zeros_like(probs)
    56 target_one_hot[np.arange(num_samples), labels] = 1
    57
    58 pred_cdf = np.cumsum(probs, axis=1)
    59 target_cdf = np.cumsum(target_one_hot, axis=1)
    60
    61 rps = np.mean(np.sum((pred_cdf - target_cdf)**2, axis=1))
    62 return ’RPS’, rps
    63
    64import xgboost as xgb
    65
    66# 1. Convert training data
    67dtrain = xgb.DMatrix(X_train, label=y_train_labels)
    68
    69# 2. Parameters
    70params = {
    71 ’objective’: ’rank:pairwise’,
@@
    83}
    84
    85# 3. Train model
    86num_boost_round = 100
    87xgb_rps_model = xgb.train(
@@ 9.1.3 Model Performance Comparison
The performance is evaluated using Mean Squared Error (MSE), Mean Absolute Error (MAE), and the Ranked Probability Score (CRPS).
@@ 9.1.4 Summary
  • RPS Advantage: Both the PyTorch and XGBoost models trained with the custom CRPS objective provide superior probabilistic forecasts.
@@
  • Conclusion: While MSE and MAE (as specific scoring functions, see Gneiting and Ranjan (2011)) provide insights into point prediction, CRPS is the appropriate metric for evaluating complex conditional distributions. The XGBoost model with the custom CRPS objective demonstrated the most robust performance.

### 13 Mar 2026 v2

@@ 6 Outlook
Proper scoring rules provide a principled framework for evaluation and should be reported in benchmarks on tabular regression. However, several open questions remain for tabular foundation models: Gneiting pointed out in Gneiting and Ranjan (2011) that asymmetric risk settings (finance, weather forecasting, etc.) may require optimizing for specific point functionals via tailored scoring rules. A fundamental tension arises between:
  • a general-purpose foundation model that is not modified after pretraining, and
  • the requirement that the optimal scoring rule be use-case-specific and therefore supplied by the user.
Two potential routes to resolve this tension are: (i) fine-tuning on a user-supplied scoring rule at deployment time, or (ii) task-token conditioning, where the model is prompted with a desired scoring rule or parameter of some scoring rule family for which it was trained to produce optimal forecasts under that rule. The second approach faces a fundamental challenge: there exist infinitely many proper scoring rules Gneiting and Raftery (2007), spanning infinitely many parametric families, so it is unclear whether pretraining with task tokens can generalize to arbitrary scoring objectives. High-stakes applications (finance, pharmaceutical production) with asymmetric risk may therefore require model variants pretrained with carefully chosen task-specific conditioning or fine-tuning.
Finally, a structural limitation of all scoring-rule-based approaches is that they are inherently ill-suited for eliciting properties realized only in rare or tail events Lerch et al. (2017); Brehmer and Strokorb (2019). Where training data is absent (e.g. in tail regions), no scoring rule can compensate; this is fundamentally a problem of epistemic uncertainty. PFNs trained to optimize expected scoring rules should therefore be applied with care when risk is concentrated in the tails or outside the support of the supplied context.
@@ 8 Appendix
8.1 Toy Models to investigate finite sample effects
@@ 8.1.2 Toy Models
Probabilistic Classification Models These models predict probabilities across $N=50$ discrete bins of the $y$-variable. The neural networks are optimized using the Adam optimizer and are simple multi-layer perceptrons (MLPs) featuring an input layer, two hidden layers with 64 neurons each using $\tanh$ activations, and a Softmax output layer for $n$ bins.
@@
Loss Function CRPS Loss CrossEntropyLoss
@@
  • PyTorch CRPS Model: A neural network trained directly to minimize the expected CRPS, which optimizes the accuracy of the cumulative distribution function (CDF).
  • XGBoost CRPS Model: An XGBoost vector-leaf classifier that uses a custom objective function to directly optimize the Continuous Ranked Probability Score:
@@
    7 Improved CRPS Objective for XGBoost.
@@
    46import xgboost as xgb
    47
    48# 1. Convert training data
    49dtrain = xgb.DMatrix(X_train, label=y_train_labels)
    50
    51# 2. Parameters
    52params = {
    53 ’objective’: ’rank:pairwise’, #overwritten later by custom objective
@@
    65}
    66
    67# 3. Train model
    68num_boost_round = 100
    69xgb_crps_model = xgb.train(
@@ 8.1.3 Model Performance Comparison
The performance is evaluated using Mean Squared Error (MSE), Mean Absolute Error (MAE), and the Continuous Ranked Probability Score (CRPS).
@@ 8.1.4 Summary
  • CRPS Advantage: Both the PyTorch and XGBoost models trained with the custom CRPS objective provide superior probabilistic forecasts.
@@
  • Conclusion: While MSE and MAE (as specific scoring functions, see Gneiting and Ranjan (2011)) provide insights into point prediction, CRPS is the appropriate metric for evaluating complex conditional distributions.
8.2 Detailed Results
8.3 Beta(1.8)-Energy Score Fine-Tuning realTabPFNv2.5
We find improved performance in terms of $R^{2}$:
Table 6: Comparative Performance Metrics (Fine-Tuned vs. Base). W/L/T counts the number of datasets where the fine-tuned model exceeded the baseline by more than a small $\varepsilon$ (Win), the baseline was better by more than $\varepsilon$ (Loss), or the absolute difference was at most $\varepsilon$ (Tie).
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +4.28% $\pm$ 10.06% +1.46% 61/21/8
RMSE +2.16% $\pm$ 4.39% +0.80% 61/24/5
$R^{2}$ +1.47pp $\pm$ 3.73pp +0.25pp 52/19/19
CRPS +2.76% $\pm$ 6.38% +0.79% 61/23/6
IntervalScore95 +2.24% $\pm$ 13.86% +0.94% 55/33/2
Table 7: Improvement by Dataset: Fine-Tuned vs. Base
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRPS Imp. IS95 Imp.
Abalone -0.03% +0.47% +0.39pp +0.78% +0.60%
Ailerons -0.04% -0.24% -0.06pp +0.18% -0.30%
Bike Sharing +1.49% +1.11% +0.13pp +1.08% -0.49%
Black Friday +0.16% +0.30% +0.30pp +0.25% -0.74%
Brazilian Houses -7.33% +3.43% +6.28pp +0.99% +5.64%
California Housing +0.91% +1.10% +0.48pp +1.25% +1.37%
Cpu Small +1.38% -0.12% -0.00pp -0.03% +0.92%
Diamonds +0.61% +1.26% +0.05pp +0.53% -1.46%
Digits +34.54% +6.29% +0.01pp +8.27% +13.74%
Elevators +0.02% +0.20% +0.03pp +0.21% +0.82%
House 16H -0.39% -0.06% -0.08pp -0.47% -0.69%
Kin8nm +1.83% +2.05% +0.30pp +2.01% +5.18%
Mercedes Benz +15.71% +6.94% +7.33pp +13.58% +7.27%
NYC Taxi +8.64% +4.73% +6.68pp +8.18% +3.50%
Protein +1.44% +0.26% +0.25pp +0.51% +1.12%
Pol +8.31% +3.30% +3.71pp +5.75% +9.80%
Puma8NH +7.87% +7.95% +0.77pp +6.85% +7.06%
Sulfur +1.85% -0.04% -0.02pp -0.17% -13.03%
Figure 8: Impact of beta=1.8 energy score fine-tuning on realTabPFNv2.5
8.4 CRLS Fine-Tuning realTabPFNv2.5
Table 8: Overall Performance Improvement (Fine-Tuned vs. Base)
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +0.65% $\pm$ 10.13% +0.21% 54/41/5
RMSE +1.67% $\pm$ 4.60% +0.96% 73/25/2
$R^{2}$ +0.94pp $\pm$ 2.40pp +0.33pp 64/20/16
CRLS +2.27% $\pm$ 7.00% +1.47% 79/18/3
IntervalScore95 +3.87% $\pm$ 13.15% +1.76% 75/22/3
Table 9: Detailed Dataset Improvements: Fine-Tuned vs. Base
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRLS Imp. IS95 Imp.
Abalone -0.76% +0.42% +0.35pp +1.52% +2.58%
Ailerons +1.01% +0.65% +0.17pp +1.52% +0.86%
Ames Housing -0.21% +1.41% -0.23pp -0.75% -7.59%
Bike Sharing +1.86% +1.88% +0.21pp +1.19% +1.59%
Black Friday +0.23% +0.61% +0.61pp +1.09% +0.67%
Brazilian Houses -24.19% +3.67% +6.72pp -2.45% +23.84%
California Housing +0.92% +1.94% +0.74pp +2.48% +3.10%
Cpu Small +0.69% +0.05% +0.00pp +1.14% +1.45%
Diamonds +0.60% +1.27% +0.05pp +1.01% +1.15%
Digits +28.23% +7.48% -0.02pp +10.91% +3.09%
Elevators -0.54% +0.13% +0.02pp +0.34% +0.67%
House 16H -1.88% +0.42% +0.27pp +0.20% +0.61%
House Sales -0.71% -1.35% -0.34pp +0.22% -1.02%
Kin8nm -0.39% -0.77% -0.12pp -0.32% +0.97%
Mercedes Benz +3.46% +1.14% +1.22pp +2.12% +1.31%
NYC Taxi +4.71% +3.04% +3.90pp +4.19% +2.08%
Protein -1.47% +0.51% +0.45pp +1.43% +3.34%
Pol +7.37% +3.23% +3.64pp +8.15% +13.82%
Puma8NH +7.24% +7.41% +0.71pp +6.26% +6.37%
Sulfur -13.24% +0.36% +0.48pp +5.14% +18.50%
Figure 9: Impact of CRLS fine-tuning on realTabPFNv2.5
8.5 CRLS performances of TabICLv2 versus realTabPFNv2.5
Table 10: Aggregate Performance Improvement (TabICLv2 vs. realTabPFNv2.5)
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +1.41% $\pm$ 12.65% +0.16% 56/48/1
RMSE +1.85% $\pm$ 10.02% +0.96% 65/34/6
$R^{2}$ +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
CRLS +6.01% $\pm$ 10.95% +3.49% 84/19/2
IntervalScore95 +5.14% $\pm$ 14.18% +2.15% 70/35/0
Table 11: Dataset-Specific Improvement Metrics
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRLS Imp. IS95 Imp.
Abalone +0.22% +0.92% +0.76pp +3.37% +4.94%
Ailerons -1.69% -1.01% -0.29pp +1.55% +0.23%
Ames Housing +0.37% +8.13% +32.38pp +4.95% -13.51%
Bike Sharing +0.21% -0.62% -0.06pp +2.83% -0.13%
Black Friday +0.59% +1.17% +1.18pp +0.34% -2.20%
Brazilian Houses -13.82% +3.41% +6.13pp +25.02% +31.86%
California Housing +4.87% +5.25% +1.91pp +8.33% +6.34%
Cpu Small +1.49% +0.08% +0.00pp +6.17% +3.52%
Diamonds +1.15% +1.62% +0.06pp +4.33% +4.93%
Digits +42.92% +21.15% +0.89pp +34.26% +43.52%
Elevators -0.20% +0.57% +0.10pp +1.55% -0.55%
House 16H +0.76% +2.92% +1.92pp +5.30% +5.42%
House Sales -5.00% -10.86% -2.37pp +0.11% -3.01%
Kin8nm +0.59% -0.29% -0.04pp +1.52% +3.26%
Mercedes Benz +18.83% +9.57% +9.92pp +20.31% +13.27%
NYC Taxi -7.26% +3.99% +9.25pp -5.47% -6.48%
Protein +2.63% +1.71% +1.55pp +4.28% +3.15%
Pol +10.39% +4.05% +4.56pp +11.22% +7.74%
Puma8NH -13.29% -13.86% -1.50pp -13.66% -10.89%
Sulfur -10.13% +1.27% +1.61pp +6.38% +16.39%
Visualizing Soil -4.08% -0.36% -0.32pp +3.50% +0.08%
Figure 10: Impact of changing from realTabPFNv2.5 to TabICLv2: Even stronger outperformance of TabICLv2 over realTabPFNv2.5 on CRLS than that achieved by fine-tuning realTabPFNv2.5 on CRLS
8.6 CRPS performances of realTabPFNv2.5 and TabICLv2
Table 12: Overall Model Improvement (TabICLv2 vs. realTabPFNv2.5)
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +1.41% $\pm$ 12.65% +0.16% 56/48/1
RMSE +1.85% $\pm$ 10.02% +0.96% 65/34/6
$R^{2}$ +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
CRPS +2.01% $\pm$ 10.17% +0.73% 59/45/1
IntervalScore95 +5.14% $\pm$ 14.18% +2.15% 70/35/0
Table 13: Dataset-Specific Performance Gains
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRPS Imp. IS95 Imp.
Abalone +0.22% +0.92% +0.76pp +0.86% +4.94%
Ailerons -1.69% -1.01% -0.29pp -2.20% +0.23%
Ames Housing +0.37% +8.13% +32.38pp -1.81% -13.51%
Bike Sharing +0.21% -0.62% -0.06pp -0.75% -0.13%
Black Friday +0.59% +1.17% +1.18pp +0.61% -2.20%
Brazilian Houses -13.82% +3.41% +6.13pp +3.63% +31.86%
California Housing +4.87% +5.25% +1.91pp +5.84% +6.34%
Cpu Small +1.49% +0.08% +0.00pp +1.82% +3.52%
Diamonds +1.15% +1.62% +0.06pp +1.45% +4.93%
Digits +42.92% +21.15% +0.89pp +28.19% +43.52%
Elevators -0.20% +0.57% +0.10pp -0.17% -0.55%
House 16H +0.76% +2.92% +1.92pp +1.82% +5.42%
House Sales -5.00% -10.86% -2.37pp -4.79% -3.01%
Kin8nm +0.59% -0.29% -0.04pp +0.20% +3.26%
Mercedes Benz +18.83% +9.57% +9.92pp +20.91% +13.27%
NYC Taxi -7.26% +3.99% +9.25pp -8.96% -6.48%
Protein +2.63% +1.71% +1.55pp +2.79% +3.15%
Pol +10.39% +4.05% +4.56pp +6.73% +7.74%
Puma8NH -13.29% -13.86% -1.50pp -15.15% -10.89%
Sulfur -10.13% +1.27% +1.61pp +1.78% +16.39%
Visualizing Soil -4.08% -0.36% -0.32pp -0.56% +0.08%